OpenStudy (stephanos100):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <10, 0>, v = <0, -9>

OpenStudy (zpupster):

check your dot product a x b if dot product = 0 they are orthogonal if they are equal they are parallel u x v = (10)(0) + (0)(9)= 0+0=0

OpenStudy (stephanos100):

so orthogonal then?

OpenStudy (loser66):

yup

OpenStudy (stephanos100):

thanks